OpenStudy (anonymous):

150=1000e^-k30 thanks for reply but more help needed- can you show all the steps!

OpenStudy (anonymous):

\[150=\frac{ 1000 }{ e ^{30k} }\] k=? Right?

OpenStudy (anonymous):

\[\frac{150}{1000}=e^{-30k}\] \[ \ln \frac{150}{1000} = -30k\]

OpenStudy (anonymous):

yep got there -can you keep going

OpenStudy (anonymous):

just divide by -30 ^_^

OpenStudy (anonymous):

you could also do this \[\ln \Bigg[ \Big(\frac{15}{100}\Big)^{-\frac{1}{30}}\Bigg]=k\] but it's equivalent to \[k=-\frac{1}{30}\ln\Big(\frac{15}{100}\Big)\]

OpenStudy (anonymous):

sorry im struggling In 0.15 =k-30 -1.897=k-30 so k =0.06323 except book says thats the wrong answer ??

OpenStudy (anonymous):

what answer does the book give?

OpenStudy (anonymous):

k=0.005417

OpenStudy (anonymous):

hmm... i sec...

OpenStudy (anonymous):

I think the book is wrong... you can get that number with 150=1000e^(-350.2k) But other than that I dunno. Wolfram says that it's wrong, too (hit the approximate form on the real solution) http://www.wolframalpha.com/input/?i=150%3D1000e^%28-30x%29

OpenStudy (anonymous):

thanks !!! hope the teacher agrees